OpenStudy (anonymous):

the density of osmium, which is the densest metal, is 22.57 g/cm^3. What is the mass of a block of osmium that measures 1.00 cm x 2.50 cm? show your work

OpenStudy (cwrw238):

density = mass / volume so mass = volume * density

OpenStudy (cwrw238):

you give measures of the block as 1.00 * 2.50 which is an area not a volume

OpenStudy (anonymous):

what is the mass of a block of osmium that measures 1.00cm x 4.00 cm x 2.50 cm?

OpenStudy (anonymous):

is that what you need?

OpenStudy (cwrw238):

yes - so plug this value of the volume and the density into the formula for the mass

OpenStudy (anonymous):

ok so would it be 22.57 g / 10cm^3?

OpenStudy (cwrw238):

ok - ill check

OpenStudy (cwrw238):

no its 22.57 * 10 g

OpenStudy (cwrw238):

volume * density

OpenStudy (anonymous):

so i multiply them?

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

'*' = multiply

OpenStudy (anonymous):

oh ... sorry didnt realize that

OpenStudy (cwrw238):

thats ok

OpenStudy (anonymous):

so would the answer be .02257 grams? I'm kind of confused...

OpenStudy (cwrw238):

22.57 * 10 = 225.7 g when multiplying by 10 move decimal point one position to the right

OpenStudy (anonymous):

oh ok that helps. so i moved it the wrong direction

OpenStudy (cwrw238):

examples 34.5 * 10 = 345 by 100 , 2 to the right: 23.67 * 100 = 2367 when dividing you move decimal to the left eg 56.89 / 10 = 5.689

OpenStudy (anonymous):

alright, thanks so much for your help!

OpenStudy (cwrw238):

yw